OpenStudy (strawberry17):

Find the exact value of the trig function: cos(-2π/3)

OpenStudy (john_es):

You must remember that \[\cos(\pi/3)=1/2\\ \sin(\pi/3)=\sqrt{3}/2\] and use, \[\cos(2\pi/3)=\cos^2(\pi/3)-\sin^2(\pi/3)\] Could you continue from this point?

OpenStudy (strawberry17):

I wish :O no, please help further... :/

OpenStudy (zzr0ck3r):

\[\cos(a) = \cos(-a)\]

OpenStudy (strawberry17):

So, cos(-2π/3) is the same thing as cos(2π/3)?

OpenStudy (zzr0ck3r):

correct

OpenStudy (strawberry17):

So then it would be -1/2 ...? :)

OpenStudy (zzr0ck3r):

correct

OpenStudy (strawberry17):

okay that makes much more sense! Thanks : )

OpenStudy (strawberry17):

who should I give the medal to? you both helped really well ...

OpenStudy (zzr0ck3r):

the other guy

OpenStudy (strawberry17):

ok Lol